OpenStudy (anonymous):

Convert each of the following to degrees: 7pi/4, 11pi/3 Convert each of the following to degrees: 7pi/4, 11pi/3 @Mathematics

OpenStudy (anonymous):

multiply by 180degrees/pi

OpenStudy (anonymous):

315degrees and 660degrees

OpenStudy (anonymous):

To change from radians to degrees you multiply the value in radians by \[\frac{180}{\pi}\] So the first one would be \[\frac{7\pi}{4} * \frac{180}{\pi}\] The pi's cancel out leaving you with \[\frac{7*180}{4}\] Which is \[315\]